OpenStudy (anonymous):

How many grams of aluminum chloride must decompose to produce 78.3 milliliters of aluminum metal, if the density of aluminum is 2.70 g/mL?

OpenStudy (jfraser):

use the density of aluminum as a conversion factor (or fraction)

OpenStudy (jfraser):

you need to do this in a multi-step process. You need to start with the volume of aluminum you need to produce, then find the mass of aluminum that creates it. you need to balance the reaction first. Can you do that?

OpenStudy (jfraser):

\(\_AlCl_3 \rightarrow \_Al + \_Cl_2\)

OpenStudy (jfraser):

it's important that you use a reaction arrow, not an equals sign. chemical reactions are not the same thing as algebraic equations

OpenStudy (jfraser):

you need to produce 78.3mL of aluminum, but the coefficients of a balanced reaction don't use mL or even grams, they use MOLES, right?

OpenStudy (jfraser):

you need to take the mL of aluminum the problem gives you, and turn it into MOLES you need to use a conversion fraction to do that

OpenStudy (jfraser):

the density of aluminum is a great fraction, because it has both MASS and VOLUME in its unit (g/mL)

OpenStudy (jfraser):

\[78.3\cancel{mL \space Al} * (\frac{2.7g \space Al}{1\cancel{mL \space Al}})\]

OpenStudy (jfraser):

cancelling the UNITS in a problem will always work better than trying to remember whether or not to multiply or divide

OpenStudy (jfraser):

yes, and what UNIT does it have?

OpenStudy (jfraser):

exactly, GJ

OpenStudy (jfraser):

that's CLOSER to moles, but we can't use the equation yet

OpenStudy (jfraser):

do you have a relationship between the grams of a substance and its moles?

OpenStudy (jfraser):

the relationship we use as the conversion fraction is the MOLAR MASS

OpenStudy (jfraser):

yes

OpenStudy (jfraser):

i'm going to use 27 just because i'm a little lazy, but see how the molar mass has 2 units in it, and it's in a fraction?

OpenStudy (jfraser):

\[211.41\cancel{g Al} * (\frac{1mol \space Al}{27\cancel{g \space Al}})\]

OpenStudy (jfraser):

good. NOW we can use the balanced reaction.

OpenStudy (jfraser):

See how 2 moles of \(Al\) are produced every time 2 moles of \(AlCl_3\) are used?

OpenStudy (jfraser):

the coefficients in the balanced reaction act just like another conversion fraction

OpenStudy (jfraser):

\[7.83\cancel{mol \space Al} * (\frac{2mol \space AlCl_3}{2\cancel{mol \space Al}})\]

OpenStudy (jfraser):

you divided by 2, but you have to multiply by 2 again, look at the fraction, it's 2/2

OpenStudy (jfraser):

it is, but now it has units of MOLES of \(AlCl_3\), not moles of \(Al\)

OpenStudy (jfraser):

almost. there's 1 more step

OpenStudy (jfraser):

do you see how we used the molar mass to go from grams of Al to moles of Al?

OpenStudy (jfraser):

well we're going to do the same thing, but upside down, to go from moles back to grams

OpenStudy (jfraser):

but we need the molar mass of \(AlCl_3\)

OpenStudy (jfraser):

good. now use that as the last fraction

OpenStudy (jfraser):

\[7.83\cancel{mol \space AlCl_3}* (\frac{133.3g \space AlCl_3}{1\cancel{mol \space AlCl_3}})\]

OpenStudy (jfraser):

that's it

OpenStudy (jfraser):

YW. stoichiometry problems are basically all the same. Find the moles of what you start with, use the coefficients from the balanced reaction to convert, and then find the grams of the "other"